Give an indirect proof of the following statement.
If 3x-2 = 7, then x ≠ 4

Answer by Edwin McCravy(20064)

If 3x-2 = 7, then x ≠ 4

For contradiction, assume that 3x - 2 = 7 yet x = 4

Then we can substitute 4 for x in

3x - 2 = 7

So therefore

3(4) - 2 = 7

  12 - 2 = 7

      10 = 7

But 10 does not equal to 7.  Therefore we have 
reached a contradiction, which proves that the 
assumption

3x - 2 = 7 and x = 4 is false. 

Therefore if 3x - 2 = 7, then x ≠ 4 

Edwin